# Entropy Rectifying Guidance for Diffusion and Flow Models

### v1

@@ 4.3 Analysis and ablations
Text Attention Energy.  To isolate the effect of the temperature re-scaling in the conditioning, we experiment with C-ERG and disable temperature rescaling in the image denoising model. In Figure 4, we vary the SoftMax temperature $\tau_{c}$ used in the text encoder, and consider generation performance for different guidance strengths $\omega$. For all metrics we find a somewhat symmetrical behavior around $\tau_{c}=1$, which corresponds to an unguided prediction because in this case there is no difference with the normal conditional prediction. The CLIPScore, coverage, density, and precision metrics generally improve when moving away from $\tau_{c}=1$, making it either larger or smaller, yet the reverse is seen for the recall metric. For FID the best values are obtained for intermediate temperature values for any setting of the guidance weight $\omega$. While lower temperatures yield the best CLIPSCore, higher terperatures are best for coverage and FID. Compared with standard classifier-free guidance (left-hand side of the dashed lines), we find that all metrics can be improved for any guidance scale, provided with the right temperature.
@@
Combining the different parts.  In Table 5, we combine different parts of the Entropy Rectifying Guidance and measure the effects on different facets of image generation. Our results show that all components show positive effects on the different facets of generation. Most of the improvement in CLIPScore are brought by the conditional entropy rectification (C-ERG, second column), while the improvements in Coverage and Density mostly come from modifying the image denoiser to obtain ERG (last column).
@@
Multistep gradient descent.  In Table 5 we the effect of varying the step-size $\gamma$ and number of updates $K$ for each attention operation. We find small variations in metrics with respect to the baseline $K=1,\gamma=1$ with slight positive variations when setting $\gamma=1.5$, using multiple gradient descent steps did not induce significant gains in our case. Therefore, we used $K=1,\gamma=1$ in our default setup in our experiments, unless specified otherwise.
Table 4: Multi-step gradient descent for optimizing the energy landscape. The baseline corresponds to $K=1,\tau=1$.
$K$ $1$ $1$ $10$
$\gamma$ $1.5$ $0.5$ $0.1$
$\Delta$ FID $0.0$ $+0.56$ $+0.94$
$\Delta$ Precision $+1.02$ $+1.44$ $+5.37$
$\Delta$ Recall $+0.36$ $-1.8$ $-0.02$
$\Delta$ Density $+3.4$ $+1.7$ $-2.88$
$\Delta$ Coverage $+0.86$ $-0.58$ $-1.1$
$\Delta$ CLIP $-0.053$ $-0.068$ $-0.09$
$\Delta$ VQA $+0.71$ $-0.95$ $-0.0058$
@@
Table 5: Impact of the different components of ERG. We successively add different components of ERG, mainly C-ERG through conditioner entropy rectification, then I-ERG through denoiser entropy rectification and objective reweighting, finally all components are merged into ERG.
Conditioner Entropy $\tau_{c}$
Denoiser Entropy $\tau_{i}$
Energy Reweighting $\alpha$
Density $98.24$ $109.52$ $120.25$ ${\bf 123.65}$
Coverage $71.12$ $72.06$ $73.21$ ${\bf 74.07}$
CLIPScore $26.45$ $26.73$ ${\bf 26.86}$ $26.81$
@@ 5 Conclusion
We presented a novel guidance mechanism for sampling diffusion and flow models – Entropy Rectifying Guidance (ERG) – which significantly improves quality without sacrificing diversity and consistency performance. In particular, by manipulating the energy landscape of the attention layers in the diffusion transformer and the text encoder at inference time, ERG significantly boosts the performance of the different models when studying their quality-consistency-diversity trade-offs and is applicable to different modalities such as text-to-image, class conditional and unconditional models. Furthermore, ERG outperforms different baselines such as CADS, APG, SEG and AutoGuidance while requiring the same compute as standard classifier-free guidance. ERG can be seamlessly combined with approaches such as APG and CADS in order to further improve sampling.
@@ Appendix A Definitions and Background
Background on Diffusion.  Consider a data point ${\bf x}$ drawn from the distribution $p_{data}(x)$, and let $t$ be a time step within the interval $[0,1]$. The forward process is defined as ${\bf x}_{t}={\bf x}+\sigma(t)\boldsymbol{\epsilon}$, where Gaussian noise $\boldsymbol{\epsilon}\sim N(0,\boldsymbol{I})$ is added to the data. The function $\sigma(t)$ is monotonically increasing, with boundary conditions $\sigma(0)=0$ and $\sigma(1)=\sigma_{\text{max}}$, where $\sigma_{\text{max}}$ is significantly larger than $\sigma_{\text{data}}$. Karras et al. (2022) demonstrated that the progression of the noisy samples ${\bf x}_{t}$ can be captured by an ordinary differential equation (ODE):
$d{\bf x}=-\dot{\sigma}(t)\sigma(t)\nabla_{{\bf x}_{t}}\log p_{t}({\bf x}_{t})dt$ (11)
@@
$\begin{split}d{\bf x}=&-\dot{\sigma}(t)\sigma(t)\nabla_{{\bf x}_{t}}\log p_{t}({\bf x}_{t})dt\\ &-\beta(t)\sigma(t)^{2}\nabla_{{\bf x}_{t}}\log p_{t}({\bf x}_{t})+\sqrt{2\beta(t)\sigma(t)}d\boldsymbol{\omega}_{t}\end{split}$ (12)
In this context, $d\omega_{t}$ represents the standard Wiener process, and $p_{t}({\bf x}_{t})$ denotes the distribution of the perturbed samples, with initial and final conditions $p_{0}=p_{data}$ and $p_{1}=\mathcal{N}(0,\sigma^{2}_{max}I)$, respectively.
To sample from diffusion models, one must solve the diffusion ODE or SDE in reverse, moving from $t=1$ back to $t=0$. This process relies on the time-dependent score function $\nabla_{{\bf x}_{t}}\log p_{t}({\bf x}_{t})$, which is approximated by a denoiser $v_{\theta}({\bf x}_{t},t)$. This denoiser is trained to predict the original clean samples ${\bf x}$ from their noisy counterparts ${\bf x}_{t}$.
The framework also supports conditional generation by employing a denoiser $v_{\theta}({\bf x},t,y)$ that incorporates additional input signals $y$, such as class labels or textual prompts, allowing for more controlled and specific sample generation.
Latent Diffusion Models (LDMs).  Are a class of generative models, proposed by Rombach et al. (2022), that operate in a latent space induced by a pretrained and frozen autoencoder. The autoencoder converts all images into latents of smaller resolution, inducing a space in which the generative model is trained, afterwards the latents can be converted back into image space by using the decoder of the autoencoder.
A.1 Rank Scoring
@@
Appendix B Additional Experiments
B.1 C-ERG Impact on diversity
@@
To better understand the effect of c-ERG on diversity, we perform a simple experiment where we track the initial velocity direction during sampling and how it varies when the initial noised input ${\bf x}_{0}=\boldsymbol{\epsilon}\sim\mathcal{N}(0,I)$ varies.
@@
We examine the variance of the predicted velocity at the start of sampling, conditioning either on the encoding of an empty prompt $\boldsymbol{\phi}_{\emptyset}$ or on the encoding of caption using temperature rescaled attention layers $\boldsymbol{\phi}^{\tau}_{c}$. For this we sample $N=20$ random noise inputs and compare the variance when the inputs vary, the right panel of Figure 5 presents a histogram of the variances where each datapoint corresponds to different spatial/channel location; we refer to it this as marginal variance:
@@
$\begin{cases}\mathrm{Var}_{\begin{subarray}{c}{\epsilon\sim\mathcal{N}(0,I)}\\ {c\sim D_{c}}\end{subarray}}\left[v_{\Theta}(\boldsymbol{\epsilon},\boldsymbol{\phi}_{\emptyset},t=0)\right],\\ \mathrm{Var}_{\begin{subarray}{c}{\epsilon\sim\mathcal{N}(0,I)}\\ {c\sim D_{c}}\end{subarray}}\left[v_{\Theta}(\boldsymbol{\epsilon},\boldsymbol{\phi}^{\tau}_{c},t=0).\right]\end{cases}$ (13)
@@
Similarly, in the right panel of Figure 5 we consider the variance of the difference between these velocity estimates and the true conditional one, i.e. when we condition on $\boldsymbol{\phi}_{c}$; we refer to this as conditional variance:
@@
From Figure 5, we observe that standard classifier-free guidance results in a larger range for the conditional variance while the marginal variance is significantly reduced. In contrast, C-ERG exhibits a larger range of marginal variances but a smaller range of conditional variances. Such results indicate that standard classifier-free guidance with high guidance scales operate as an initial condition for the flow ODE (equation 11) that has low variance, resulting in exploring only a subset of the solutions of the ODE. Furthermore, as the initial conditional velocity is much larger, it can easily cause overshooting problems as large updates are being taken with a flawed prediction because of imperfections in the model, while the low marginal variance results in oversaturated colors and simplified image compositions. Similar observations were made by Saharia et al. (2022); Kynkäänniemi et al. (2024); Sadat et al. (2025).
Qualitative results corroborate this, see for example Figure 1 where using standard CFG results in low diversity, saturated images with simple compositions, while our ERG is able to generate more complex scenes that suffer less from the “cartoonish” effect, especially for highly out-of-distribution prompts where the denoiser model is expected to be less accurate.
B.2 I-ERG Impact on quality
To better understand the differences between I-ERG and CFG, we consider the difference between the two denoising terms used in CFG and I-ERG, see Equation 14. In particular we decompose the difference into parallel and orthogonal components to the conditional prediction, and track the magnitude of these components throughout the sampling process.
Results in Figure 6 reveal a fundamental difference between I-ERG and CFG. The parallel differences show similar trends between CFG and I-ERG: a peak around $t=0.075$ followed by a sharp decline towards zero, indicating that for approximately $t\geq 0.4$ the denoising task becomes easy enough that both conditional and unconditional predictions yield similar and correlated results. On the other hand, when examining the orthogonal component of the differences, we see that it keeps a steady norm through sampling while in the case of standard CFG, the norm of the orthogonal difference quickly decreases as sampling progresses for $t\geq 0.4$, this indicates strong correlations between conditional and unconditional predictions as the denoising task becomes easier and the denoiser needs less to rely on the conditional embeddings.
While Sadat et al. (2025) completely eliminate the parallel component from the difference term, we are able to break such correlations by manipulating the energy landscape of the attention layers of the denoiser.
Imposing this divergence between the positive and negative predictors results in semantically meaningful errors for the negative velocity prediction which, when used as a guidance term, improves the low-level details in the resulting image.

### v2

@@ 5.3 Analysis and ablations
Text attention energy.  To isolate the effect of the temperature re-scaling in the text-encoder and image denoising components, we experiment with C-ERG and disable temperature rescaling in the image denoising model. In Figure˜4, we vary the SoftMax temperature $\tau_{c}$ used in the text encoder, and consider generation performance for different guidance strengths $\omega$. For all metrics we find a somewhat symmetrical behavior around $\tau_{c}=1$, which corresponds to an unguided prediction because in this case there is no difference with the normal conditional prediction.
The CLIPScore, Coverage and Density metrics generally improve when moving away from $\tau_{c}=1$, making it either larger or smaller. For FID, the best values are obtained with low guidance scales and intermediate temperature values, the general trend shows improved FID using C-ERG compared to standard guidance. Compared with standard classifier-free guidance (left-hand side of the dashed lines), we find that all metrics can be improved for any guidance scale, provided with the right temperature. While $\tau_{c}<1$ results in higher CLIPScore, $\tau_{c}>1$ results in higher Coverage, indicating that tuning $\tau_{c}$ provides an easy way to control the diversity-consistency tradeoff.
@@
Combining the different parts.  In Table˜5, we combine different parts of ERG and measure the effects on different facets of image generation. Our results show that all components show positive effects across all metrics, at the expense of a slight degradation in FID. Compared to the CFG baseline (first row), most of the improvement in CLIPScore and VQA are brought by the conditional entropy rectification ($\tau_{c}$, second row), while the improvements in Coverage and Density mostly come from rectifying the attention in the denoiser to obtain ($\tau_{i}$, third row). Finally, a further modest improvement in Density, Coverage and VQA is brought by the update step size ($\gamma$, fourth row).
@@
Multi-step gradient descent.  In Table˜5 we consider the effect of varying number of updates $K$ for each attention operation along with the update step size. We find small variations in metrics with respect to the baseline $K=\gamma=1$ with slight improvements when setting $\gamma=1.5$, using multiple gradient descent steps did not induce significant gains. Therefore, we used $K=\gamma=1$ in our default setup in our experiments, unless specified otherwise.
@@
Table 4: Impact of the different components of ERG. We accumulate different components of ERG, namely C-ERG, then I-ERG through denoiser entropy rectification and objective reweighting, finally all are merged into ERG.
$\tau_{c}$ $\tau_{i}$ $\gamma$ FID $(\downarrow)$ Density $(\uparrow)$ Coverage $(\uparrow)$ CLIP $(\uparrow)$ VQA $(\uparrow)$
12.81 98.24 71.12 26.45 70.15
13.06 109.52 72.06 26.73 73.10
[…]
13.62 123.65 74.07 26.81 74.67
Table 5: Multi-step gradient descent for optimizing the energy landscape. we ablate different values for $K$ and $\gamma$.
[…]
@@ 6 Conclusion
We presented Entropy Rectifying Guidance (ERG), a novel guidance mechanism for sampling diffusion and flow models which significantly improves sample quality without sacrificing diversity and consistency performance. In particular, by manipulating the energy landscape of the attention layers in the diffusion transformer and the text encoder at inference time, ERG significantly boosts the performance of different models when studying their quality-consistency-diversity trade-offs and is applicable to different modalities such as text-to-image, class conditional and unconditional models. Furthermore, ERG outperforms recent state-of-the-art guidance methods such as CADS, APG, SEG and AutoGuidance, while requiring the same compute as standard CFG. ERG can be combined with approaches such as APG and CADS, which further improves results.
Acknowledgements.  Karteek Alahari was supported in part by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korean Government (MSIT) (No. RS-2024-00457882, National AI Research Lab Project).
@@ Appendix A Definitions and Background
A.1 Preliminaries
Background on diffusion.  Consider a data point $\bf x$} drawn from the distribution $p_{data}({\bf x})$, and let $t$ be a time step within the interval $[0,1]$. The forward process is defined as ${\bf x}_{t}={\bf x}+\sigma(t)\boldsymbol{\epsilon}$, where Gaussian noise $\boldsymbol{\epsilon}\sim N(0,\boldsymbol{I})$ is added to the data. The function $\sigma(t)$ is monotonically increasing, with boundary conditions $\sigma(0)=0$ and $\sigma(1)=\sigma_{\text{max}}$, where $\sigma_{\text{max}}$ is significantly larger than $\sigma_{\text{data}}$. Karras et al. (2022) demonstrated that the progression of the noisy samples ${\bf x}_{t}$ can be captured by an ordinary differential equation (ODE):
$d{\bf x}=-\dot{\sigma}(t)\sigma(t)\nabla_{{\bf x}_{t}}\log p_{t}({\bf x}_{t})dt.$ (5)
@@ A.1 Preliminaries
$\begin{split}d{\bf x}=&-\dot{\sigma}(t)\sigma(t)\nabla_{{\bf x}_{t}}\log p_{t}({\bf x}_{t})dt\\ &-\beta(t)\sigma(t)^{2}\nabla_{{\bf x}_{t}}\log p_{t}({\bf x}_{t})+\sqrt{2\beta(t)\sigma(t)}d\boldsymbol{\omega}_{t}.\end{split}$ (6)
In this context, $d\boldsymbol{\omega}_{t}$ represents the standard Wiener process, and $p_{t}({\bf x}_{t})$ denotes the distribution of the perturbed samples, with initial and final conditions $p_{0}=p_{data}$ and $p_{1}=\mathcal{N}(0,\sigma^{2}_{max}I)$, respectively.
To sample from diffusion models, one must solve the diffusion ODE or SDE in reverse, moving from $t=1$ back to $t=0$. This process relies on the time-dependent score function $\nabla_{{\bf x}_{t}}\log p_{t}({\bf x}_{t})$, which is approximated by a denoiser $v_{\theta}({\bf x}_{t},t)$. This denoiser is trained to predict the original clean samples $\bf x$} from their noisy counterparts ${\bf x}_{t}$. The framework also supports conditional generation by employing a denoiser $v_{\theta}({\bf x},t,y)$ that incorporates additional input signals $y$, such as class labels or textual prompts, allowing for more controlled and specific sample generation.
Background on flow matching.  Flow matching approaches aim to learn a velocity field $v_{t}$ mapping random noise ${\bf x}_{0}=\boldsymbol{\epsilon}\sim\mathcal{N}(0,I)$ to data samples ${\bf x}_{1}\sim p_{\text{data}}({\bf x})$. Such a mapping is obtained by solving an ordinary differential equation (ODE) of the form:
$\frac{d{\bf x}_{t}}{dt}=v_{t}({\bf x}_{t}).$ (7)
Lipman et al. (2023) provide a simple simulation-free training objective for flow generative models by directly regressing the velocity field $v_{t}$ on a conditional vector field $u_{t}(.|{\bf x}_{1})$ :
$\mathbb{E}_{t,q(x_{1}),p_{t}(x_{t}|x_{1})}\lVert v_{t}({\bf x}_{t})-{\bf u}_{t}({\bf x}_{t}|{\bf x}_{1})\rVert^{2},$ (8)
where ${\bf u}_{t}(.|{\bf x}_{1})$ uniquely determines a conditional probability path $p_{t}(.|{\bf x}_{1})$ towards data sample ${\bf x}_{1}$. A popular choice for the conditional probability path corresponds to a linear interpolation between data and noise $\forall t\in[0,1],\ {\bf x}_{t}=t{\bf x}_{1}+(1-t){\bf x}_{0}$ resulting in a conditional probability path of the form ${\bf u}({\bf x}_{t}|{\bf x}_{1})={\bf x}_{1}-{\bf x}_{0}$. Once the conditional probability path is learned, sampling from the model can be achieved by solving the ODE defined in Equation˜7 using any appropriate ODE solver from the literature.
Latent diffusion models (LDMs).Rombach et al. (2022) proposed to train diffusion models in a latent space induced by a pretrained and frozen autoencoder. The autoencoder converts all images into latents of smaller resolution, inducing a space in which the generative model is trained, afterwards the latents can be converted back into image space by using the decoder of the autoencoder. Such models allow to significantly scale up model training as the effective number of tokens representing each data point is reduced significantly. For example, recent autoencoders have a downscaling factor of $8\times 8$, resulting in a reduction of the sequence length by a factor of 64.
A.2 Rank scoring
@@
Appendix B Limitations
While our method is successfully able to boost the Pareto fronts for quality-consistency-diversity, tradeoffs between these metrics still subsist, although to a lesser extent, when using ERG. Future work should explore whether these tradeoffs are inherent to the model or a result of subpar sampling methods.
Additionally, our method operates under the assumption that the model follows the diffusion transformer architecture. As our method manipulates the energy landscape in attention layers, it needs to be tuned for different architectures containing a different number of attention layers. It is therefore not directly applicable to models that do not make use of attention layers in their architecture, such as U-Nets without self-attention.
Our method was only tested on diffusion transformers with standard architecture shapes (presenting a depth between $16$ and $38$), therefore extrapolating our claims to extremely deep/shallow models is not evident without supporting experiments. However, we believe such cases to be less relevant currently as they are not usual choices in state-of-the-art generative models.
Appendix C Interpreting ERG
In the following, we provide two interpretations of ERG based on entropy regularized RL and variational inference. Let $p$ be the density function for the strong model and $p^{\tau}$ the one induced by the attention rectification mechanism we introduce in the paper.
I-ERG.  We begin with the I-ERG guidance update, which modifies the model’s score estimate as:
$\nabla_{z}\log p^{\tau}({\bf z})=\nabla_{z}\log p({\bf z})\;+\;w\,\big(\nabla_{z}\log\rho({\bf z})-\nabla_{z}\log p^{\tau}({\bf z})\big).$ (9)
Rewriting the right-hand side, we obtain:
$\nabla_{z}\log p^{\tau}({\bf z})=\nabla_{z}\log p({\bf z})\;+\;w\,\nabla_{z}\log\!\left(\frac{\rho({\bf z})}{p^{\tau}({\bf z})}\right).$ (10)
Integrating both sides with respect to $\bf z$}, we find $p^{\tau}({\bf x})\propto p({\bf x})\cdot\exp\big(w\cdot R({\bf x})\big)$, where $R({\bf x})=\log\!\left(\frac{\rho({\bf x})}{p({\bf x})}\right)$.
This form reveals that I-ERG guidance parallels a posterior distribution that re-weights the base model’s density $p({\bf x})$ by an exponential function of the reward signal $R({\bf x})$. This exponential reweighting aligns with objectives used in entropy-regularized RL or KL control, where the optimal distribution maximizes an expected reward under a KL penalty with respect to a prior:
$\pi^{\star}=\arg\max_{\pi}\,\mathbb{E}_{x\sim\pi}[R({\bf x})]-\frac{1}{\lambda}\mathrm{KL}(\pi||p),$ (11)
where $\lambda$ serves as an inverse temperature controlling the sharpness of reward influence. The solution to this optimization problem has the form:
$\pi({\bf x})\propto p({\bf x})\cdot\exp\big(\lambda R({\bf x})\big).$ (12)
This shows that the I-ERG update implicitly implements the maximum entropy policy framework, steering the sampling distribution toward high-reward regions while maintaining proximity to the base model $p({\bf z})$. Equivalently, using Bayes rule, it can be viewed as defining a joint distribution over $\bf x$} under $p({\bf x})$ and the energy induced by reward $R({\bf x})$ (assuming conditional independence).
C-ERG.  In the conditional setting, assume we start from a model $p({\bf z}|{\bf s})$ conditioned on some semantic signal $\bf s$}. Let ${\bf s}^{\tau}={\bf s}+\Delta{\bf s}$ represent a degraded or weakened version of the conditioning (e.g., via temperature scaling or perturbation). The C-ERG guidance update is given by:
$\nabla_{z}\log p^{\mathrm{ERG}}({\bf z}\mid{\bf s},{\bf s}^{\tau})=\nabla_{z}\log p({\bf z}\mid{\bf s})+w\cdot\big(\nabla_{z}\log p({\bf z}\mid{\bf s})-\nabla_{z}\log p({\bf z}\mid s^{\tau})\big).$ (13)
Or, equivalently:
$\nabla_{z}\log p^{\mathrm{ERG}}({\bf z}\mid{\bf s},{\bf s}^{\tau})=(1+w)\,\nabla_{z}\log p({\bf z}\mid{\bf s})-w\,\nabla_{z}\log p({\bf z}\mid{\bf s}^{\tau}).$ (14)
To analyze this further, we use a first-order Taylor expansion of the degraded score $\nabla_{z}\log p({\bf z}\mid{\bf s}^{\tau})$ around $\bf s$}:
$\nabla_{z}\log p({\bf z}\mid{\bf s}^{\tau})\approx\nabla_{z}\log p({\bf z}\mid{\bf s})+\frac{\partial}{\partial{\bf s}}\big(\nabla_{z}\log p({\bf z}\mid{\bf s})\big)\cdot\Delta{\bf s}.$ (15)
Substituting this into the update:
$\nabla_{z}\log p^{\mathrm{ERG}}({\bf z}\mid{\bf s},{\bf s}^{\tau})\approx\nabla_{z}\log p({\bf z}\mid{\bf s})-w\cdot\left(\frac{\partial}{\partial{\bf s}}\big(\nabla_{z}\log p({\bf z}\mid{\bf s})\big)\cdot\Delta{\bf s}\right).$
Thus, C-ERG introduces a Jacobian-level product correction, nudging the sample along directions where the score is most sensitive to perturbations in the conditioning. This helps reinforce fine-grained semantic detail in generation, even when the conditioning signal is noisy or coarsened.
This Jacobian-product reward has a natural interpretation: it reflects how the image-level score $\nabla_{z}\log p({\bf z}\mid{\bf s})$ changes when the conditioning $\bf s$} is slightly perturbed. This resembles classifier guidance, where one steers the generation based on $\nabla_{z}\log p({\bf y}\mid{\bf z})$—the gradient of a classifier with respect to the image.
Here, the conditioning $\bf s$} is a continuous vector, and C-ERG guides generation using:
$\frac{\partial}{\partial s}\big(\nabla_{z}\log p({\bf z}\mid{\bf s})\big)\cdot\Delta{\bf s},$ (16)
which is the direction in image space most sensitive to changes in the semantics of $\bf s$}. This amounts to amplifying features that are especially responsive to semantic precision.
Appendix D Additional Experiments
D.1 C-ERG Impact on diversity
@@
To better understand the effect of c-ERG on diversity, we perform a simple experiment where we track the initial velocity direction during sampling and how it varies when the initial noised input ${\bf x}_{1}=\boldsymbol{\epsilon}\sim\mathcal{N}(0,I)$ varies.
@@
We examine the variance of the predicted velocity at the start of sampling, conditioning either on the encoding of an empty prompt $\boldsymbol{\phi}_{\emptyset}$ or on the encoding of caption using temperature rescaled attention layers $\boldsymbol{\phi}^{\tau}_{c}$. For this we sample $N=20$ random noise inputs and compare the variance when the inputs vary, the right panel of Figure˜5 presents a histogram of the variances where each datapoint corresponds to different spatial/channel location; we refer to it this as marginal variance:
@@
$\begin{cases}\mathrm{Var}_{\begin{subarray}{c}{\epsilon\sim\mathcal{N}(0,I)}\\ {c\sim D_{c}}\end{subarray}}\left[v_{\Theta}(\boldsymbol{\epsilon},\boldsymbol{\phi}_{\emptyset},t=0)\right],\\ \mathrm{Var}_{\begin{subarray}{c}{\epsilon\sim\mathcal{N}(0,I)}\\ {c\sim D_{c}}\end{subarray}}\left[v_{\Theta}(\boldsymbol{\epsilon},\boldsymbol{\phi}^{\tau}_{c},t=0)\right],\end{cases}$ (17)
@@
Similarly, in the left panel of Figure˜5 we consider the variance of the difference between these velocity estimates and the true conditional one, i.e. when we condition on $\boldsymbol{\phi}_{c}$; we refer to this as conditional variance:
@@
From Figure˜5, we observe that standard classifier-free guidance results in a larger range for the conditional variance while the marginal variance is significantly reduced. In contrast, C-ERG exhibits a larger range of marginal variances but a smaller range of conditional variances. Such results indicate that standard classifier-free guidance with high guidance scales operate as an initial condition for the flow ODE (equation˜7) that has low variance, resulting in exploring only a subset of the solutions of the ODE. Furthermore, as the initial conditional velocity is much larger, it can easily cause overshooting problems as large updates are being taken with a flawed prediction because of imperfections in the model, while the low marginal variance results in oversaturated colors and simplified image compositions. Similar observations were made by Saharia et al. (2022); Kynkäänniemi et al. (2024); Sadat et al. (2025).
Qualitative results corroborate this, see for example Figure˜1 where using standard CFG results in low diversity, saturated images with simple compositions, while our ERG is able to generate more complex scenes that suffer less from the “cartoonish” effect, especially for highly out-of-distribution prompts where the denoiser model is expected to be less accurate.
D.2 I-ERG impact on quality
To better understand the differences between I-ERG and CFG, we consider the difference between the two denoising terms used in CFG and I-ERG, see Equation˜18. In particular we decompose the difference into parallel and orthogonal components to the conditional prediction, and track the magnitude of these components throughout the sampling process.
Results in Figure˜6 reveal a fundamental difference between I-ERG and CFG. The parallel differences show similar trends between CFG and I-ERG: a peak around $t=0.075$ followed by a sharp decline towards zero, indicating that for approximately $t\geq 0.4$ the denoising task becomes easy enough that both conditional and unconditional predictions yield similar and correlated results. On the other hand, when examining the orthogonal component of the differences, we see that it keeps a steady norm through sampling while in the case of standard CFG, the norm of the orthogonal difference quickly decreases as sampling progresses for $t\geq 0.4$, this indicates strong correlations between conditional and unconditional predictions as the denoising task becomes easier and the denoiser needs less to rely on the conditional embeddings.
While Sadat et al. (2025) completely eliminate the parallel component from the difference term, we are able to break such correlations by manipulating the energy landscape of the attention layers of the denoiser. Imposing this divergence between the positive and negative predictors results in semantically meaningful errors for the negative velocity prediction which, when used as a guidance term, improves the low-level details in the resulting image.